OpenStudy (anonymous):

Medal for answer The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples! (written response)

OpenStudy (anonymous):

square and add, i.e. check that \[7^2+24^2=25^2\] and that \[9^2+40^2=41^2\]

OpenStudy (anonymous):

49+576=625 81+1600=1681